It takes a groundskeeper 40 minutes to prepare a Little League baseball field for a game. It takes his assistant 55 minutes to prepare the same field. How long will it take if they work together to prepare the field?
Answer by fcabanski(1391)

Remember distance problems? The formula is Distance = rate x time. Work problems have a similar formula: Work = rate x time. Work is often 1 as in 1 complete job.


The groundskeeper's rate is 1=r*40 so r = 1/40 of the job per minute. The assistant works at a rate of 1/55 of the job per minute (1 = r*55 so r=1/55).


To find the time when they work together, use the equation with their combined rate: 1/40 + 1/55 = 19/440


1 = 19/440 * t so t = 440/19 = 23 and 3/19 minutes or 23 minutes 9.47 seconds.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)